What is the negative solution of :
(3x-7)^2
thanks

Answer by unlockmath(1122)

x^2-2x+6=2x^2-6x-26
Hello,
Let's put everything on one side so it equals zero.
0 = x^2 - 4x - 32 (now we can factor this.)
0 = (x-8)(x+4)
x=8
x=-4 (here's the negative solution)
RJ toftness
www.math-unlock.com